OpenStudy (anonymous):

A circle with circumference of 10 has area of 100. True or False

OpenStudy (anonymous):

Circumference of a circle = \(2\pi r\) Area of a circle = \(\pi r^{2}\) so you can solve for \(r\) from the circumference, then solve for the area with the equation to see if it was true or false!

OpenStudy (anonymous):

Would it be 78.5?

OpenStudy (anonymous):

no. \(r\) of the circle =\(\frac{10}{2\pi}\) and area of the circle = \(\pi (\frac{10}{2\pi})^2=\frac{100}{4\pi}\) so it's not 100!

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

no problem!